# Towards Scalable IoT Deployment for Visual Anomaly Detection via Efficient Compression

### v1

@@ 6 Conclusions and future directions
In this work, we proposed a modular and efficient visual anomaly detection (VAD) framework tailored for IoT scenarios, where resource constraints at the edge pose significant challenges. By investigating both image and feature compression strategies, we demonstrated that it is possible to significantly reduce communication overhead while preserving high anomaly detection accuracy. Our approach, validated on the MVTec AD dataset using a lightweight CNN and PatchCore as the detection model, achieves performance comparable to uncompressed pipelines, with reduced latency and memory requirements, making it suitable for practical industrial deployments.
Future research will explore dynamic adaptation strategies that adjust the compression method and level based on real-time constraints such as available bandwidth or computational load. Additionally, we plan to investigate learning-based feature compression techniques and extend the framework to support online model updates and real-time anomaly localization across heterogeneous edge-server systems.

### v2

@@ 5 Conclusions and future directions
We have proposed a modular and efficient visual anomaly detection (VAD) framework tailored for IoT scenarios, where edge resource constraints pose major challenges. By investigating both image and feature compression strategies, we show that communication overhead can be significantly reduced while maintaining high detection accuracy. Validated on the MVTec AD dataset using a lightweight CNN and PatchCore, our approach matches the performance of uncompressed pipelines with lower latency and memory usage, making it viable for industrial deployment.
Future work will explore dynamic adaptation strategies that adjust compression based on real-time factors like bandwidth and computational load. We also plan to investigate learning-based feature compression and extend the framework to support online model updates and real-time anomaly localization across heterogeneous edge-server systems.